A fruit stand sells two varieties of strawberries: standard and deluxe. A box of standard strawberries sells for $3, and a box of deluxe strawberries sells for $6. In one day the stand sold 150 boxes of strawberries for a total of $645. How many boxes of each type were sold?
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

A fruit stand sells two varieties of strawberries: standard and deluxe. A box of standard strawberries sells for $3, and a box of deluxe strawberries sells for $6. In one day the stand sold 150 boxes of strawberries for a total of $645. How many boxes of each type were sold?
----------------
p = $3 boxes
3p + 6(150-p) = 645
3p + 900 - 6p = 645
-3p = -255
p = 85
===========
3*85 + 6*65 = 645

Answer by greenestamps(13200)



A quick informal solution method....

If all 150 boxes were standard, the cost would have been $450.
The actual total cost was $645, which is $195 more.
The difference in price between the two kinds of boxes is $3.
So the number of deluxe boxes sold was 195/3 = 65.

ANSWER: 65 deluxe boxes; 150-65 = 85 standard boxes